# Ordering kinetics in the active Ising model

### v1

@@ IV.2 Dynamical scaling
Figure 3: (Color online) (a–c) Scaled correlation functions $C(r/R,t)$ versus $r/R$, for the evolution of the $2d$ AIM shown in Fig. 2. Unscaled versions are shown in the inset. “Liquid” and “coexistence” signify the region where the system is quenched, $\epsilon$ denotes particle speed. (d) Log-log plot of $1-C(r/R)$ versus $r/R$. The cusp exponent is estimated as $\alpha\sim 1$. Parameters correspond to Fig. 2.
@@
In the theory of phase-ordering kinetics, the scaling hypothesis states that if the system is characterized by a single length scale $R(t)$ [$R(t)$ is equivalent to the average domain size], the domain morphology is statistically the same at all times, apart from a scale factor. When all domain lengths are measured in units of $R(t)$, the equal-time pair correlation function should exhibit the dynamical-scaling property of Eq. (6). The equal-time correlation function is a non-equilibrium quantity so as $R(t)$ which can be estimated from the decay of the correlation function. To identify the self-similar behavior of the evolving domains we plot the correlation function in Fig. 3. The scaled $C(r/R,t)$ and unscaled $C(r,t)$ correlation functions are shown in Fig. 3(a–c) for $\epsilon=0$ and $\epsilon=1$. $R(t)$ is determined from the distance over which the correlation function decays to e.g., 0.2 of its maximum value, that is, $C(r,t)=0.2C(0,t)$.
@@
In Fig. 4, we plot the scaled structure factor, $S(k,t)R(t)^{-2}$ versus $kR$, which is the Fourier transform of the correlation function. In Fourier space, Eq. (10) translates into the following power-law behavior of the structure factor: $S(k)\sim k^{-(d+\alpha)}$ and therefore, the large-$k$ behavior of the structure factor tail generates a slope $\sim-3$ (in log-log plot) for $d=2$ and $\alpha=1$ which denotes “Porod’s decay”: $S(k)\sim k^{-(d+1)}$, associated with scattering from sharp interfaces Bray (2002, 1994). This naturally originates from the long-range ordering in AIM leading to compact high-density clusters with smooth boundaries. Domains with rough morphologies having fractal interfaces do not follow Porod’s law and the large-$k$ tail of the scaled structure factor yields a non-integer exponent Chatterjee et al. (2018). A similar violation of the Porod law was also observed in the coarsening of the VM due to the irregular morphology associated with the cluster boundaries Dey et al. (2012); Katyal et al. (2020). $C(r,t)$ [and consequently, $S(k,t)$] exhibits two distinct power laws for small and large $r/R$ limits [large and small $kR$ limits] in the VM Dey et al. (2012) which we do not observe in the AIM. Furthermore, density fluctuations might also play a role in determining whether a system will follow or violate Porod’s law. In the VM, giant density fluctuations break large liquid domains and restrict the formation of large compact domains (which eventually manifests in the microphase separation of the coexistence region in the steady-state) Solon et al. (2015) and might be responsible for the non-Porod behavior of the system Dey et al. (2012). On the other hand, AIM obeys the Porod behavior where density fluctuations are normal in the liquid phase Solon and Tailleur (2015) and the steady-state manifests a bulk phase separation.
@@ IV.3 Growth law
Figure 5: (Color online) (a) $R(t)\sim t^{1/2}$ for a quench into the coexistence and liquid regimes. The late time saturation of $R(t)$ signifies that the system has reached the corresponding NESS. As the density is higher for the liquid quench, the saturation appears late. Parameters: $\rho_{0}=3$ (coexistence), $\rho=10$ (liquid), $L=400$, $\beta=1.1$ and $\epsilon=1$. (b) $R(t)$ versus $t$ for different thermal noises ($T=\beta^{-1}$, coexistence regime). Parameters: $L=400$, $\rho_{0}=3$ and $\epsilon=1$. (c) Role of self-propulsion on the characteristic length $R(t)$ as a function of $t$. For a larger $\epsilon$, the system reaches the NESS faster while for $\epsilon=0$, the corresponding dynamics is slow. Parameters: $L=400$, $\beta=1.1$, and $\rho_{0}=5$. The growth law $R(t)\sim t^{1/2}$ is unaffected by thermal fluctuations and particle activity.
@@
Therefore, the immediate question one can ask is whether the LCA growth law also governs the late-stage coarsening dynamics of an active system with non-conserved scalar order parameters such as the AIM and whether the coarsening process is diffusion-dominated. To characterize this, we plot the length scale data with time in Fig. 5 for various control parameters and quench regimes. We find that the late-stage growth kinetics of the domain in the coexistence and liquid regimes exhibit a $R(t)\sim t^{1/2}$ growth law [Fig. 5(a)]. We extract the same growth law for different temperatures [Fig. 5(b)], and self-propulsion velocities [Fig. 5(c)]. Domains identified by the correlation of density and magnetization fields also show similar growth behavior (see Appendix A). An interesting feature of the domain growth kinetics is that the domain size for a quench into the liquid regime is larger than the corresponding domain size in the coexistence regime [see Fig. 2 at $t=10^{3}$ and Fig.5(a)], although the domain morphologies for both these quench regimes are self-similar. The larger domains in the liquid regime are likely to be arising from higher density. Also, notice that a larger $\beta$ in Fig. 5(b) signifies a reduced thermal noise that slows down the local ordering of spins. Thus smaller $\beta$ allows the formation of larger domains. Nevertheless, different thermal fluctuations exhibit the same growth as thermal noise is asymptotically irrelevant for ordering in systems that are free from disorder Puri (2009).
@@
Fig. 5(c) shows the increasing length scale with time for different $\epsilon$. For $\epsilon=0$, the system is purely diffusive, and domain growth proceeds via the coarsening of connected domains (curvature-driven growth facilitated by diffusing domain wall) [see Fig. 2(c–d)]. The domain morphology looks very similar to the evolution of an Ising ferromagnet quenched below the critical temperature. Thus, a $\sim t^{1/2}$ growth law for $\epsilon=0$ similar to the pure Ising model is physically plausible. But, when $\epsilon>0$, the domains of each spin no longer remain connected and form high-density clusters that self-propel along the horizontal direction. As time progresses, these clusters spread in the transverse direction due to constant diffusion ($D=1$) and merge with other clusters. Therefore, the domain growth for $\epsilon>0$ is again a diffusive phenomenon, and consequently, the growth kinetics exhibits a $R(t)\sim t^{1/2}$ growth law similar to $\epsilon=0$. We have identified this novel mechanism of diffusion-driven domain growth in the AIM after a thorough investigation of the altered diffusion coefficient in Sec. IV.4 and Appendix B.
@@
It should also be noted that while approaching the NESS via coarsening, the high-density AIM domains of individual spins, upon merging, try to minimize the surface energy by decreasing the surface area similar to the dynamics of the passive non-conserved scalar order parameter. Therefore, although the system is active, the critical mechanisms (diffusion-dominated growth and minimization of surface energy during coarsening) of domain coarsening in AIM are similar to the $2d$ Ising model, and thus, it is not surprising that we extract a $R(t)\sim t^{1/2}$ growth law for both the passive and active models. For a more precise quantification of the asymptotic growth law, we determine the effective growth exponent, defined as:
@@
Figure 6: (Color online) Effective growth exponent $Z_{\rm eff}$ versus time $t$ for (a) different self-propulsion ($\epsilon=0,0.5$, and $1$, $\beta=1.1$ and $\rho_{0}=3$) and (b) quenching into different NESS ($\rho_{0}=3$ for coexistence and $\rho_{0}=10$ for liquid, $\beta=1.1$ and $\epsilon=1$). The dashed lines are a guide to the eyes.
@@ IV.4 Role of transverse diffusion
Figure 8: (Color online) $R(t)$ versus $t$ (on a log-log scale) showing a 0.5 growth exponent for different transverse diffusion $D_{\perp}=0.3$ and $D_{\perp}=1$. Inset: $R(t)$ versus $t$ in the absence of any transverse diffusion ($D_{\perp}=0$) showing no growth of domains with time. Parameters: $L=400$, $\beta=1.1$, $\epsilon=1$, and $\rho_{0}=8$.
To quantify the role of the magnitude of transverse diffusion coefficient ($D_{\perp}$), we plot $R(t)$ versus $t$ in Fig. 8 for two different values of $D_{\perp}$. The plot shows that the system exhibits the power-law growth of $R(t)\sim t^{1/2}$ for different values of $D_{\perp}$. Interestingly, even a small $D_{\perp}=0.3$ is sufficient to drive the proper domain growth. A larger $D_{\perp}$ only increases the average domain size. As discussed earlier in Fig. 7(b–c), a vanishing $D_{\perp}$ can not initiate a domain growth as $R(t)$ remains constant with $t$ (see inset of Fig. 8). Therefore, transverse diffusion plays the most crucial role in the AIM growth kinetics.
@@ IV.5 Growth law from the hydrodynamic description of the AIM
In this section, we want to investigate whether the continuous description of the AIM also manifests the same time dependency $(\sim t^{1/2})$ of the coarsening length scale as observed in our numerical analysis. We consider refined mean-field equations similar to Ref. Solon and Tailleur (2015) for the spatiotemporal evolutions of the density $(\rho)$ and magnetization $(m)$ fields:
@@
where $v=2D\epsilon$, $\alpha=4\beta^{2}(1-2\beta/3)$, and $r=3\alpha\alpha_{m}/2$ is a positive function of $\beta$ Solon and Tailleur (2015). In constructing Eq. (13), we slightly modify the flipping rate equation of Eq. (2) to read $W_{\rm flip}=\exp\left(-2\beta\sigma m/\rho\right)$. For the mean-field equations, $r=0$, which can not capture the AIM physics correctly as the system always exhibits homogeneous profiles (gas and liquid), and the inhomogeneous phase-separated profiles are never observed Solon and Tailleur (2015). In Eqs. (12) and (13), local fluctuations are taken into account, which are generally neglected in the mean-field approximations Solon and Tailleur (2015).
@@
Eqs. (12) and (13) immediately show the importance of the transverse diffusion. If we start with a $x$-independent initial condition, for instance, a horizontal thin stripe of width $w$, where: $\rho(y)=\rho_{0}/w$, $m(y)=m_{0}>0$ is the stationary solution of Eq. (13) for $0<y<w$, and $\rho(y)=\delta\ll\rho_{0}/w$ and $m_{0}=0$ otherwise. Then the solution stays $x$-independent at all later times: $\rho(x,y,t)=\tilde{\rho}(y,t)$ and $m(x,y,t)=\tilde{m}(y,t)$ where the equations for $\tilde{\rho}$ and $\tilde{m}$ are:
@@
Eq. (14) shows that the thin stripe expands diffusively in the $y$-direction (transverse direction) yielding a $\sqrt{t}$ dependence of the stripe width. For general initial conditions, we use explicit Euler FTCS (Forward Time Centered Space) Press (2007) differencing scheme to numerically integrate Eqs. (12) and (13). We solve these two coupled partial differential equations on a square domain of size $L\times L$ with periodic boundary conditions applied in both directions. In our simulation, $L=400$ and the maximum simulation time is $t_{\rm sim}=5\times 10^{6}$. To maintain the numerical stability criteria, we set $\Delta x=1$ as the discretization in space and $\Delta t=10^{-3}$ as the discretization in time. These discretization parameters satisfy the Courant-Friedrichs-Lewy (CFL) stability condition. In our numerical implementation, we fix $D=r=v=1$ and the initial system is prepared as a high-noise homogeneous gas phase with $\rho=\rho_{0}$ and $m=0$ by adding a zero-mean scalar Gaussian white noise to Eq. (13Solon et al. (2015). We then calculate the spatial dependence of the density and magnetization correlation using Eqs. (4) and (5) for 25 independent realizations. Finally, $R(t)$ is determined where the ensemble-averaged correlation functions decay to 0.2 of its maximum value.
@@
In Fig. 9, we plot the time evolution of the magnetization field $m$ by solving Eqs. (12) and (13) via the finite difference FTCS scheme. The formation of self-propelling clusters with smooth interfaces and their growth with time resembles the dynamics shown in Fig. 2 for the time-evolution of the microscopic model [Eqs. (13)]. We also extract the LCA growth law $R(t)\sim t^{1/2}$ (as shown in Fig. 10 on a logarithmic scale) by solving the hydrodynamic equations after quenching the system from a disordered gaseous phase to an ordered liquid phase. The length scale in Fig. 10 is obtained from the equal-time spatial correlation of the density fields although the length scale obtained from the correlation of the magnetization fields also exhibits the same growth law. Therefore, the self-propulsion terms in the hydrodynamic equations (12) and (13) of the AIM do not affect the asymptotic growth law of the non-conserved Model A.

### v2

@@ IV.2 Dynamical scaling
Figure 3: (Color online) (a–c) Scaled correlation functions $C(r/R(t),t)$ versus $r/R(t)$, for the evolution of the $2d$ AIM shown in Fig. 2. Unscaled versions are shown in the inset. “Liquid” and “coexistence” signify the region where the system is quenched, $\epsilon$ denotes particle speed. (d) Log-log plot of $1-C(r/R(t))$ versus $r/R(t)$. The cusp exponent is estimated as $\alpha\sim 1$. The parameters correspond to Fig. 2.
@@
In the theory of phase-ordering kinetics, the scaling hypothesis states that if the system is characterized by a single length scale $R(t)$ [$R(t)$ is equivalent to the average domain size], the domain morphology is statistically the same at all times, apart from a scale factor. When all domain lengths are measured in units of $R(t)$, the equal-time pair correlation function should exhibit the dynamical scaling property of Eq. (8). The equal-time correlation function is a non-equilibrium quantity so as $R(t)$ which can be estimated from the decay of the correlation function. To identify the self-similar behavior of the evolving domains we plot the correlation function in Fig. 3. The scaled $C(r/R(t),t)$ and unscaled $C(r,t)$ correlation functions are shown in Fig. 3(a–c) for $\epsilon=0$ and $\epsilon=1$. $R(t)$ is determined from the distance over which the correlation function decays to e.g., 0.2 of its maximum value, that is, $C(r,t)=0.2C(0,t)$.
@@
In Fig. 4, we plot the scaled structure factor, $S(k,t)R(t)^{-2}$ versus $kR(t)$, which is the Fourier transform of the correlation function. In Fourier space, Eq. (12) translates into the following power-law behavior of the structure factor: $S(k)\sim k^{-(d+\alpha)}$ and therefore, the large-$k$ behavior of the structure factor tail generates a slope $\sim-3$ (in log-log plot) for $d=2$ and $\alpha=1$ which denotes “Porod’s decay”: $S(k)\sim k^{-(d+1)}$, associated with scattering from sharp interfaces Bray (2002, 1994). This naturally originates from the long-range ordering in AIM leading to compact high-density clusters with smooth boundaries. Domains with rough morphologies having fractal interfaces do not follow Porod’s law and the large-$k$ tail of the scaled structure factor yields a non-integer exponent Chatterjee et al. (2018). A similar violation of the Porod law was also observed in the coarsening of the VM due to the irregular morphology associated with the cluster boundaries Dey et al. (2012); Katyal et al. (2020). $C(r,t)$ [and consequently, $S(k,t)$] exhibits two distinct power laws for small and large $r/R(t)$ limits [large and small $kR(t)$ limits] in the VM Dey et al. (2012) which we do not observe in the AIM. Furthermore, density fluctuations might also play a role in determining whether a system will follow or violate Porod’s law. In the VM, giant density fluctuations break large liquid domains and restrict the formation of large compact domains (which eventually manifest in the microphase separation of the coexistence region in the steady-state) Solon et al. (2015a) and might be responsible for the non-Porod behavior of the system Dey et al. (2012). On the other hand, AIM obeys the Porod behavior where density fluctuations are normal in the liquid phase Solon and Tailleur (2015) and the steady-state manifests a bulk phase separation.
@@ IV.3 Growth law
Figure 5: (Color online) (a) $R(t)\sim t^{1/2}$ for a quench into the coexistence (solid red square) and liquid (solid green diamond) regimes. The late time saturation of $R(t)$ signifies that the system has reached the corresponding NESS. As the density is higher for the liquid quench, the saturation appears late. Parameters: $\rho_{0}=3$ (coexistence), $\rho=10$ (liquid), $L=400$, $\beta=1.1$ and $\epsilon=1$. (b) $R(t)$ versus $t$ for different thermal noises [$T=\beta^{-1}$ with $\beta=0.9$ (solid pink circle), $1.1$ (solid red triangle) and $1.4$ (solid green inverted triangle)] in the coexistence regime. Parameters: $L=400$, $\rho_{0}=3$ and $\epsilon=1$. (c) Role of self-propulsion on the characteristic length $R(t)$ as a function of $t$. For a larger $\epsilon$ [$=0.5$ (solid green diamond) and $1.0$ (solid pink inverted pentagon)], the system reaches the NESS faster while for $\epsilon=0$ (red square), the corresponding dynamics is slow. Parameters: $L=400$, $\beta=1.1$, and $\rho_{0}=5$. The growth law $R(t)\sim t^{1/2}$ is unaffected by thermal fluctuations and particle activity.
@@
However, as outlined in the Introduction, AIM dynamics is similar to the Model C dynamics Hohenberg and Halperin (1977) where a non-conserved magnetization field is coupled to a conserved density field. In Model C, the growth law for a quench into the order-disorder phase coexisting region appears to be $R(t)\sim t^{1/3}$ (reminiscent of Model B’s growth law with a conserved order parameter Puri (2009)). Conversely, for a quench into the ordered region, the growth law follows $R(t)\sim t^{1/2}$ Kockelkoren and Chaté (2002) (reflecting the growth law of Model A with a non-conserved scalar order parameter).
@@
Therefore, the immediate question we ask is whether the late-stage growth of the coarsening length scale in the AIM (an active system) follows the Model C growth laws, which vary depending on the quenching regime. To characterize this, we plot the length scale data with time in Fig. 5 for various control parameters and quench regimes. We find that the late-stage growth kinetics of the domains in the coexistence and liquid regimes exhibit a $R(t)\sim t^{1/2}$ growth law [Fig. 5(a)]. We extract the same growth law for different temperatures [Fig. 5(b)], and self-propulsion velocities [Fig. 5(c)]. Domains identified by the correlation of density and magnetization fields also show similar growth behavior (see Appendix A). The domains exhibit larger sizes when quenching into the liquid regime compared to those formed when quenching into the coexistence region, primarily due to higher density. However, despite this difference, the coarsening length scale follows the same growth pattern in both scenarios [see Fig. 2 at $t=10^{3}$ and Fig.5(a)]. Also, notice that a larger $\beta$ in Fig. 5(b) signifies a reduced thermal noise that slows down the local ordering of spins. Thus smaller $\beta$ allows the formation of larger domains. Nevertheless, different thermal fluctuations exhibit the same growth as thermal noise is asymptotically irrelevant for ordering in systems that are free from disorder Puri (2009).
Fig. 5(c) shows the increasing length scale with time for different self-propulsion velocity $\epsilon$. For the non-motile AIM ($\epsilon=0$), the system is purely diffusive, and domain growth proceeds through the coarsening of connected domains of $\sigma=\pm 1$ (curvature-driven growth facilitated by diffusing domain wall) [see Fig. 2(c–d)]. Therefore the domain morphology of the magnetization field looks very similar to the evolution of an Ising ferromagnet quenched below the critical temperature. Thus, a $\sim t^{1/2}$ growth law for $\epsilon=0$ similar to the pure Ising model is physically plausible. Furthermore, in the case of the non-motile AIM, we observe the same growth exponent $\theta=1/2$ for the conserved density field. We hypothesize that this similarity arises from the fact that the $\epsilon=0$ limit of the AIM belongs to the same universality class as the Ising model Solon and Tailleur (2015). It is important to note that while AIM and Model C share the same combination of a non-conserved magnetization with Ising symmetry coupled to a conserved density, the crucial distinction lies in the fulfillment of detailed balance. Model C satisfies detailed balance, whereas even at $\epsilon=0$, AIM is not in equilibrium and fails to satisfy detailed balance regarding any distribution. Consequently, Scandolo et al. Scandolo et al. (2023) argued that the critical point of AIM belongs to a distinct universality class from that of Model C. Moreover, it was suggested Scandolo et al. (2023) that if a general AIM model is formulated where detailed balance gets restored in the zero self-propulsion limit, it would belong to the Model C universality class. This has recently been shown in a thermodynamically consistent model of the AIM Agranov et al. (2024).
@@
However, for $\epsilon>0$, the domains of each spin no longer remain connected and form high-density clusters that self-propel along the horizontal direction. As time progresses, these clusters spread in the transverse direction due to constant diffusion ($D=1$) and merge with other clusters. Therefore, the domain growth for $\epsilon>0$ is again a diffusive phenomenon, and consequently, the growth kinetics exhibits a growth law $R(t)\sim t^{1/2}$ similar to $\epsilon=0$. We have identified this novel mechanism of diffusion-driven domain growth in the AIM after a thorough investigation of the altered diffusion coefficient in Sec. IV.4 and Appendix B.
@@
It should also be noted that while approaching the NESS via coarsening, the high-density AIM domains of individual spins, upon merging, try to minimize the surface energy by decreasing the surface area similar to the ordering dynamics of the passive Ising model. Therefore, although the system is active, the critical mechanisms (diffusion-dominated growth and minimization of surface energy during coarsening) of domain coarsening in AIM are similar to the $2d$ Ising model, and thus, it is not surprising that we extract a $R(t)\sim t^{1/2}$ growth law for both the passive and active models. For a more precise quantification of the asymptotic growth law, we determine the effective growth exponent, defined as:
@@
Figure 6: (Color online) Effective growth exponent $Z_{\rm eff}$ versus time $t$ for (a) different self-propulsion [$\epsilon=0$ (solid red square), $0.5$ (solid blue circle), and $1$ (solid green triangle), $\beta=1.1$ and $\rho_{0}=3$] and (b) quenching into different NESS [$\rho_{0}=3$ for coexistence (solid red square) and $\rho_{0}=10$ for liquid (solid green triangle), $\beta=1.1$ and $\epsilon=1$]. The dashed lines are a guide to the eyes.
Now, we aim to highlight the differences between the coarsening dynamics in the AIM and the VM, where both systems involve the coupling of a conserved density field with a non-conserved magnetization field. The main difference between the coarsening dynamics of AIM and VM lies in their growth laws. In the VM, the coarsening length scale of the density domains grows as $R(t)\sim t^{0.25}$, while the magnetization length scale grows much faster, $R(t)\sim t^{0.83}$ Katyal et al. (2020). The slower growth of the density domains can be attributed to the fractal morphology of the density field. Although density clusters in the VM are not compact, due to activity, individual particles may move between these clusters, transmitting the orientation order over larger length scales than the sizes of the density clusters, resulting in faster growth in the coarsening length scale of the magnetization Katyal et al. (2020). On the other hand, in the AIM, both the density and magnetization fields exhibit a $R(t)\sim t^{0.5}$ growth law. The main ingredient responsible for this growth in the AIM is diffusion along the transverse direction which will be discussed in the following section. Regarding the morphology of the density and magnetization domains in these two models, the density fluctuation appears to be crucial. In the VM, density fluctuation is giant Solon et al. (2015a), whereas in the AIM, density fluctuation is normal Solon and Tailleur (2015). Giant density fluctuations break large liquid domains, arrest band coarsening, and restrict the formation of large compact domains in the VM. In contrast, we observe compact high-density clusters with smooth boundaries in the AIM. This difference elucidates why we observe a faster growth in the density domains compared to the VM. The behavior of the magnetization fluctuations in the VM and AIM mirrors that of the density fluctuations: the VM exhibits giant fluctuations, while the AIM shows normal fluctuations. Despite this difference, both VM and AIM exhibit long-range order (LRO). The LRO in AIM is a natural consequence of its discrete symmetry Solon and Tailleur (2015), while the continuous symmetry VM exhibits LRO Toner and Tu (1995) due to the non-equilibrium activeness of the particles. Therefore, the faster growth in the magnetization length scale in the VM (compared to AIM) can be attributed to activity, as discussed in Sec. IV.5.
@@ IV.4 Role of transverse diffusion
Figure 8: (Color online) $R(t)$ versus $t$ (on a log-log scale) showing a 0.5 growth exponent for different transverse diffusion $D_{\perp}=0.2$ (solid cyan diamond), 0.3 (solid yellow inverted triangle), 0.5 (solid green square), and 1 (solid blue triangle). Inset: $R(t)$ versus $t$ for small values of the transverse diffusion, $D_{\perp}=0$ (open black inverted pentagon), 0.05 (open red diamond), and 0.1 (open purple inverted triangle) showing no temporal domain growth. Parameters: $L=400$, $\beta=1.1$, $\epsilon=1$, and $\rho_{0}=8$.
To quantify the role of the diffusion coefficient ($D_{\perp}$), we plot $R(t)$ versus $t$ in Fig. 8 for various values of $D_{\perp}$. The plot shows that the system follows a power-law growth of $R(t)\sim t^{1/2}$ across different values of $D_{\perp}$. Notably, even a small value of $D_{\perp}=0.2$ is sufficient to drive proper domain growth. Increasing $D_{\perp}$ further only increases the average domain size. As shown earlier in Fig. 7(b–c), a vanishing $D_{\perp}$ or a very small $D_{\perp}$ can not initiate domain growth, as $R(t)$ remains constant with $t$ (see inset of Fig. 8). This indicates that transverse diffusion plays one of the most crucial roles in the growth kinetics of the AIM. However, our primary motivation for suppressing transverse diffusion extends beyond simply verifying the diffusion-driven nature of coarsening. As shown in Fig. 8, we explore the critical regime of transverse diffusion ($D_{\perp}$) where the growth exponent transitions to the diffusive behavior observed in Model A. By systematically reducing $D_{\perp}$ and analyzing the resulting growth dynamics, we identify the limiting values of $D_{\perp}$ at which this transition occurs. This investigation provides insights into the sensitivity of the growth kinetics of the AIM to the strength of transverse diffusion.
@@ IV.5 Growth law from the hydrodynamic description of the AIM
In this section, we want to investigate whether the continuous description of the AIM Solon and Tailleur (2013, 2015); Kourbane-Houssene et al. (2018); Scandolo et al. (2023) also manifests the same time dependency $(\sim t^{1/2})$ of the coarsening length scale as observed in our numerical analysis. We consider refined mean-field equations similar to Ref. Solon and Tailleur (2015) for the spatiotemporal evolutions of the density $(\rho)$ and magnetization $(m)$ fields:
@@
where $v=2D\epsilon$, $\alpha=4\beta^{2}(1-2\beta/3)$, and $r=3\alpha\alpha_{m}/2$ is a positive function of $\beta$ Solon and Tailleur (2015). In constructing Eq. (15), we slightly modify the flipping rate equation of Eq. (4) to read $W_{\rm flip}=\exp\left(-2\beta\sigma m/\rho\right)$. For the mean-field equations, $r=0$, which cannot correctly capture AIM physics as the system always exhibits homogeneous profiles (gas and liquid), and inhomogeneous phase-separated profiles are never observed Solon and Tailleur (2015). In Eqs. (14) and (15), local fluctuations are taken into account, which are generally neglected in the mean-field approximations Solon and Tailleur (2015).
@@
Eqs. (14) and (15) immediately show the importance of transverse diffusion. If we start with a $x$-independent initial condition, for instance, a horizontal thin stripe of width $w$, where: $\rho(y)=\rho_{0}/w$, $m(y)=m_{0}>0$ is the stationary solution of Eq. (15) for $0<y<w$, and $\rho(y)=\delta\ll\rho_{0}/w$ and $m_{0}=0$ otherwise. Then the solution stays $x$-independent at all later times: $\rho(x,y,t)=\tilde{\rho}(y,t)$ and $m(x,y,t)=\tilde{m}(y,t)$ where the equations for $\tilde{\rho}$ and $\tilde{m}$ are:
@@
Eq. (16) shows that the thin stripe expands diffusively in the $y$-direction (transverse direction) yielding a $\sqrt{t}$ dependence of the stripe width. For general random initial conditions, we use explicit Euler FTCS (Forward Time Centered Space) Press (2007) differencing scheme to numerically integrate Eqs. (14) and (15). We solve these two coupled partial differential equations on a square domain of size $L\times L$ with periodic boundary conditions applied in both directions. In our simulation, $L=400$ and the maximum simulation time is $t_{\rm sim}=5\times 10^{6}$. To maintain the numerical stability criteria, we set $\Delta x=1$ as the discretization in space and $\Delta t=10^{-3}$ as the discretization in time. These discretization parameters satisfy the Courant-Friedrichs-Lewy (CFL) stability condition. In our numerical implementation, we fix $D=r=v=1$ and the initial system is prepared as a high-noise homogeneous gas phase with $\rho=\rho_{0}$ and $m=0$ by adding a zero-mean scalar Gaussian white noise to Eq. (15Solon et al. (2015a). We then calculate the spatial dependence of the density and magnetization correlation using Eqs. (6) and (7) for 25 independent realizations. Finally, $R(t)$ is determined where the ensemble-averaged correlation functions decay to 0.2 of its maximum value.
@@
In Fig. 9, we plot the time evolution of the magnetization field $m$ by solving Eqs. (14) and (15) via the finite difference FTCS scheme. The formation of self-propelling clusters with smooth interfaces and their growth with time resembles the dynamics shown in Fig. 2 for the time-evolution of the microscopic model [Eqs. (15)]. We also extract the LCA growth law $R(t)\sim t^{1/2}$ (as shown in Fig. 10 on a logarithmic scale) by solving the hydrodynamic equations after quenching the system from a disordered gaseous phase to an ordered liquid phase. The length scale in Fig. 10 is obtained from the equal-time spatial correlation of the density fields (the length scale obtained from the correlation of the magnetization fields also exhibits the same growth law). Hence, the self-propulsion terms in the hydrodynamic equations (14) and (15) of the AIM do not alter the asymptotic growth law of the non-conserved Model A.
It can be seen that for $v=0$ $(\epsilon=0)$, Eq. (15) transforms into the deterministic version of the well-known time-dependent Ginzburg–Landau (TDGL) equation Puri (2009) of the order parameter evolution. Then one can derive the Allen-Cahn equation of motion for the interfaces from the TDGL equation Puri (2009), and this will produce a growth $R(t)\sim t^{1/2}$ as shown in Fig. 5(c). Eq. (14) for $v=0$ follows a purely diffusive equation and therefore the density length scale will also exhibit a $\sqrt{t}$ dependence.
We now aim to discuss and compare the impact of hydrodynamics on domain growth in both the AIM and the VM. From Eqs. (14) and (15), we observe that in the AIM, density is advected by the magnetization only in the $x$-direction and vice versa. Therefore, diffusion, denoted by the term $D\nabla^{2}\rho$ (or $D\nabla^{2}m$) becomes exceptionally significant in the growth of the density (magnetization) domains in the $2d$ AIM. Now, let us consider the minimal hydrodynamic equations that can describe the VM Solon et al. (2015a, b):
$\partial_{t}\rho=-v\vec{\nabla}\cdot\vec{m}\,,$ (18)
$\partial_{t}\vec{m}=D\nabla^{2}\vec{m}-\lambda\vec{\nabla}\rho-\xi\left(\vec{m}\cdot\vec{\nabla}\right)\vec{m}+\left(a_{2}-a_{4}|\vec{m}|^{2}\right)\vec{m}\,,$ (19)
where $\vec{m}(\vec{r},t)$ is the vectorial magnetization field for continuous symmetry, $a_{2}$ and $a_{4}$ are functions of $\rho$, $\lambda\vec{\nabla}\rho$ reflects the pressure gradient induced by density heterogeneities, $\xi$ and $D$ are two transport coefficients associated with the advection and the diffusion of the magnetization, respectively. As one can notice, the advective term $v\partial_{x}m$ of the AIM is replaced by $v\vec{\nabla}\cdot\vec{m}$ in Eq. (18), and in the VM, density can only be advected along the direction of the magnetization $\vec{m}$. As there is no $D\nabla^{2}\rho$ term in the Vicsek density equation, diffusion cannot play a role in the growth of the density domains by diffusing particles in directions other than those dictated by $\vec{m}$. Ordering dynamics in self-propelled particles with the diffusive term in the density equation has been reported to show a faster growth of the density field Das et al. (2018). In Eq. (19), along with the diffusive term $D\nabla^{2}\vec{m}$, the additional $\left(\vec{m}\cdot\vec{\nabla}\right)\vec{m}$ term is an advective term (analogous to the advective term in the Navier-Stokes equation). It is well known that in Model H Hohenberg and Halperin (1977), where advection is included (a conserved order parameter is coupled to hydrodynamic flow), the asymptotic growth exponent is greater than 0.5: $R(t)\sim t$ when the advective term is dominant at late times (viscous hydrodynamic regime), and even later times, when the inertial effect also becomes important (inertial hydrodynamic regime), $R(t)\sim t^{0.67}$ Bray (2003). Therefore, advection is probably most relevant for the emergence of the large value (0.83) in the growth exponent for magnetization in the VM. In contrast to the VM, within the AIM, the flow occurs exclusively along the $x$ direction, emphasizing the crucial role of diffusion in growth kinetics. As a result, both density and magnetization grow diffusively.